# Integrated Sensing and Communications Over the Years: An Evolution Perspective

### v2

@@ III-A3 Summary and Discussion
Subsequent research efforts shifted towards developing waveform designs capable of leveraging common hardware architectures and unified signal processing algorithms[9]. This stage emphasized the need for a holistic framework that not only maximizes individual link performance but also coordinates cross-domain resource utilization for global system efficiency. Nevertheless, fundamental challenges persisted, particularly in characterizing performance boundaries and in modeling the complex mechanisms of cross-domain interference. While studies such as [134] made initial progress in analyzing interference in shared-spectrum ISAC systems, this area remained underdeveloped compared to traditional communication or radar theory.
@@ III-B1 Types of Sensing Entities
UE are equipped with specialized sensors to gather data about the environment or surrounding objects and communicate with BSs using wireless signals. Owing to mobility and flexibility, multiple UEs collaboratively extended sensing coverage. For example, satellites in the global navigation satellite system (GNSS) and BSs in current cellular networks can utilize device-based sensing technology to locate mobile devices [135].
@@ III-B2 Topology of Single-Cell Network
Summary and Discussion
These limitations highlight the need for tailored strategies such as interference mitigation, coordinated beamforming, and asynchronous sensing support. A key issue is intra-cell interference arising from overlapping radar and communication functionalities at centralized BSs. In [136], a monostatic ISAC configuration was investigated, where frequency-domain null-space projection reduced interference power by more than 10 dBm while preserving sensing performance. To further improve operation, [137] proposed a coordinated multipoint framework incorporating distributed beamforming and joint SINR optimization, which achieved communication SINR gains of up to 8 dB without degrading radar detection accuracy. Beyond waveform-level coordination, [138] studied uplink sensing under asynchronous transceivers and constrained antenna resources. Cross-antenna cross-correlation accurately estimates delays and Doppler shifts without strict synchronization, effectively resolving multi-target ambiguities by exploiting spatial correlation across receive antennas.
Figure 6: Topologies of Space-air-ground Integrated Network
@@ III-B3 Topology of Multi-Cell Network
C-RAN, introduced around 2009, further advances the multi-cell paradigm by decoupling and centralizing BS functionalities. In LTE and 5G NR, a BS (gNB) can be functionally split into centralized units for higher-layer baseband processing, distributed units for lower-layer real-time tasks, and remote radio units (RRUs) for analog transmission and reception [143, 144]. To address full-duplex challenges, some RRUs are designated as dedicated receivers. When combined with massive MIMO and cooperative beamforming, C-RAN enables large-scale, coordinated wide-area sensing. Joint signal- and data-level fusion in C-RAN has been shown to improve both communication throughput and sensing accuracy [6]. In high-density networks, optimizing cluster size, user density, and spatial resource allocation to fully leverage the performance gains [145].
@@
The progression from distributed BS cooperation to heterogeneous macro–micro deployments and centralized C-RAN frameworks highlights the growing complexity of multi-cell ISAC. While these architectures offer extended coverage, spatial diversity, and improved sensing accuracy, they also introduce challenges in inter-cell synchronization, interference management, and coordination overhead. Advanced signal processing methods, such as compressed sensing and background subtraction, have been applied to recover perception parameters and mitigate multipath distortion [146]. A stochastic geometry model in [145] demonstrated that inter-cell coordination reduced sensing interference by up to 12% and improved SINR by nearly 5 dB. Moreover, joint beamforming and SINR optimization further improved worst-case radar SNR by 8% under communication constraints [147]. These findings underscore the need for unified cross-layer frameworks that can balance S&C demands in dynamic multi-cell environments.
@@ III-B4 Space-Air-Ground Integrated Network
The space–air–ground integrated network has been proposed to provide seamless connectivity and joint S&C services across multiple layers. As shown in Fig. 6, this architecture supports applications such as environmental monitoring, disaster response, and remote communications in areas where conventional infrastructure is limited. By integrating satellites, aerial platforms, and terrestrial networks, it not only extends coverage but also enables large-scale, real-time data acquisition and situational awareness across diverse geographic environments.
@@
Despite its promise, SAGIN introduces unique technical hurdles. Spectrum sharing across segments is complicated by rapidly varying interference conditions, which necessitate collaborative spectrum sensing approaches, such as the distributed cognitive satellite–ground framework proposed in [153]. Latency and real-time data processing pose additional challenges in high-mobility aerial scenarios, where distributed edge computing and predictive beamforming techniques [154] enable low-latency and mobility-resilient service continuity. Moreover, transmitting sensitive multi-segment data raises security and privacy concerns. Blockchain-based schemes have been proposed to guarantee tamper-proof, decentralized data integrity across space, air, and ground domains.
@@ III-C1 How Can Clock Asynchronization Be Effectively Mitigated in Distributed ISAC Networks
Clock asynchronization remains a fundamental limitation in large-scale ISAC deployments. Misaligned oscillators introduce timing offsets, carrier frequency shifts, and phase noise, thereby degrading range–Doppler estimation and symbol detection. Research has investigated mitigation through CSI ratio-based models [155], CRB-based performance analysis [156], and advanced estimators such as iterative cross-correlation [142], Kalman filtering [157], and optimization-based offset estimation [158]. Practical implementations, such as the SPACE framework for joint synchronization and positioning and anchor-assisted uplink sensing [159], have integrated offset estimation and compensation. Despite these advances, scalable synchronization solutions for dynamic, multi-node ISAC networks remain limited, and trade-off analyses, such as those between synchronization accuracy and resource utilization, need further development.
@@ III-C2 How Should Sensing Architectures Be Designed to Balance Latency, Fusion Accuracy, and Scalability
Efficient sensing architecture design has been recognized as a critical enabler for the evolution of ISAC networks toward multi-cell and hierarchical deployments. At the 3GPP 6G workshop in March 2025 [160], several candidate solutions were introduced, with a three-layer sensing framework emerging as one of the most intuitive options. This framework consists of three key components: (i) local sensing nodes, which include Remote Radio Units (RRUs), roadside units, and ISAC-enabled user equipment (UEs); (ii) intermediate aggregators, such as edge servers and distributed Baseband Units (BBUs); and (iii) the sensing function implemented within the core network. In this architecture, environmental signals are captured and pre-processed at the local nodes. The intermediate aggregators then collect and refine the measurements from multiple sources. Finally, the core network enforces system-wide requirements and carries out advanced data fusion and analytics. By distributing initial processing and partial fusion to these intermediate layers—rather than transmitting raw sensor data directly to the core—this hierarchical design reduces transmission latency, alleviates computational burdens on the sensing nodes, and mitigates the delays associated with fully centralized processing.
@@ IV-A From Edge Intelligence to Edge Perception
Modern smart environments are supported by billions of edge devices (e.g., smartphones, wearable sensors, and connected vehicles) that generate massive volumes of heterogeneous data. Traditional cloud-centric approaches, which offload computation to centralized servers, often incur high latency, excessive communication overhead, and elevated privacy risks [161]. Edge intelligence addresses these limitations by distributing computation across edge nodes, enabling localized data analysis and decision-making[162, 163]. Federated learning (FL), in particular, offers a privacy-preserving solution by training models locally while aggregating global insights at edge servers[164].
@@ IV-A2 Edge Perception for Sensing Enhancement
Edge perception plays a central role in enhancing sensing capabilities through multi-modal data fusion at the network edge. By employing advanced AI algorithms, subtle patterns that elude conventional methods can be captured, thereby improving detection accuracy and situational awareness. For instance, in smart homes, sensor data is monitored to assess occupant behavior and environmental factors, which in turn facilitate the automatic optimization of temperature, humidity, and lighting to enhance comfort. In industrial IoT, real-time factory monitoring is utilized to support predictive maintenance and production scheduling, while wearable devices track patient vitals to enable prompt medical interventions.
@@ IV-A3 Edge Perception for Communication Enhancement
In addition to sensing, edge perception strengthens communication performance by fusing diverse sensing modalities into the transmission process. Localized processing reduces end-to-end latency and alleviates computational pressure on centralized infrastructure. Multi-modal fusion improves communication reliability and resource management. In vehicular networks, radar, camera, and LiDAR inputs can be combined with RF signals to enhance interference mitigation, beamforming precision, and spectrum allocation. This synergy increases throughput and reduces latency. Beyond resource optimization, reinforcement learning and cross-attention architectures dynamically allocate spectrum resources, further improving adaptability and robustness in highly dynamic environments [166].
@@ IV-A4 Interplay between ISAC and AI
This synergy has been validated across multiple application domains. In vehicular networks, a transformer-based predictive beamforming scheme that jointly exploited radar and CSI data improved beam selection accuracy and link throughput under mobility, achieving up to 4.2 × gains in low-speed scenarios [170]. In indoor sensing, vision transformer–based models effectively processed Wi-Fi CSI data for human activity recognition (HAR). In particular, the class-attention in image transformers architecture achieved state-of-the-art classification accuracy on the University of Texas Human Activity Recognition (UT-HAR) dataset under constrained edge inference conditions [171]. Similarly, lightweight CSI image-based frameworks demonstrated high recognition accuracy while requiring only 2.1 MB of memory and delivering 18 frames per second throughput on Raspberry Pi-class hardware [172].
@@
This co-synergistic paradigm has been exemplified in [174], where an ISAC-guided over-the-air federated learning framework was introduced. By leveraging sensing data quality for dynamic client scheduling and power control, the framework improved convergence robustness and reduced training latency in bandwidth-limited environments. To further enhance adaptability, [25] proposed a transfer learning–based edge model adaptation framework, which enabled fine-tuning of pretrained models using few-shot local samples. This design preserved model accuracy and privacy while significantly lowering communication overhead, demonstrating the potential of ISAC-assisted adaptive learning in dynamic wireless contexts.
@@
Energy efficiency and sensing cost further represent significant barriers in resource-constrained deployments. To alleviate these constraints, a closed-loop AI-in-the-loop framework was introduced in [177], where gradient importance sampling and stochastic gradient Langevin dynamics (SGLD)-based noise modeling were applied, achieving a 52% reduction in sensing cost and a 77% saving in communication energy. In passive sensing scenarios, a deep unfolding architecture named ISAC-NET was proposed in [178], alternating between channel reconstruction and target detection. This architecture yielded a 25.6% improvement in sensing accuracy and a 31.4% reduction in channel estimation error compared with conventional methods.
@@ IV-B1 The Shift towards Task-Oriented ISAC
Figure 10: The structure in paper[179]
@@ IV-B2 Enabling AI Technologies
Among recent breakthroughs, large language models (LLMs) have demonstrated strong potential in extracting knowledge from sensory data with minimal task-specific supervision. Typically structured into three modules, including perception, grounding, and alignment, LLMs enable multimodal reasoning and semantic abstraction. For example, [180]introduced a framework where smartphone accelerometer signals were transformed into textual sequences for motion and environment inference, while electrocardiogram (ECG) signals were converted into text-based indicators such as heart rate, as shown in Fig. 10. Similarly, [179] extended LLMs beyond textual processing to encode domain knowledge into cyber-physical systems. This signal-to-text paradigm highlights the semantic abstraction capabilities of LLMs in perception-centric ISAC networks.
@@
[181] High-pass filtering on range-Doppler data CNN Gesture recognition Accuracy >96%
[182] Noise filtering and data normalization Posture detection Accuracy >93%
[183] Sparse point cloud processing CNN + BiLSTM + kNN Gesture recognition Accuracy >98%
@@
[185] Kalman filtering for motion state estimation Beam tracking in IoV systems Accuracy >92%, reduced beam misalignment by 30%
[186] Fast fourier transform (FFT) for radar signal processing, noise filtering LSTM Hand gesture recognition for contactless interfaces (vending machines, ATMs, etc.) Accuracy >92%, reduced beam misalignment by 30%
@@
[189] Cell average constant false-alarm rate algorithm, data filtering Human activity recognition (boxing, jumping, squatting, etc.) Recognition accuracy >97%, 1% higher than using single data types
[190] Miniature radar signal processing Gesture recognition Inference time <0.15 ms, Accuracy >90%
[191] Echo signal processing, spatial filtering Transformer Predictive beamforming for vehicular ISAC systems Communication sum rate >3.7 dB compared to baseline, improved performance in high mobility
@@
[193] Heartbeat-based radar sensing Supervised Learning Authentication (Heart rate data) Accuracy >97%

### v3

@@ III-A3 Summary and Discussion
Subsequent research efforts shifted towards developing waveform designs capable of leveraging common hardware architectures and unified signal processing algorithms[10418473]. This stage emphasized the need for a holistic framework that not only maximizes individual link performance but also coordinates cross-domain resource utilization for global system efficiency. Nevertheless, fundamental challenges persisted, particularly in characterizing performance boundaries and in modeling the complex mechanisms of cross-domain interference. While studies such as [liu2023fundamental] made initial progress in analyzing interference in shared-spectrum ISAC systems, this area remained underdeveloped compared to traditional communication or radar theory.
@@ III-B1 Types of Sensing Entities
UE are equipped with specialized sensors to gather data about the environment or surrounding objects and communicate with BSs using wireless signals. Owing to mobility and flexibility, multiple UEs collaboratively extended sensing coverage. For example, satellites in the global navigation satellite system (GNSS) and BSs in current cellular networks can utilize device-based sensing technology to locate mobile devices [liu2022networked].
@@ III-B2 Topology of Single-Cell Network
These limitations highlight the need for tailored strategies such as interference mitigation, coordinated beamforming, and asynchronous sensing support. A key issue is intra-cell interference arising from overlapping radar and communication functionalities at centralized BSs. In [han2024cellular], a monostatic ISAC configuration was investigated, where frequency-domain null-space projection reduced interference power by more than 10 dBm while preserving sensing performance. To further improve operation, [10577579] proposed a coordinated multipoint framework incorporating distributed beamforming and joint SINR optimization, which achieved communication SINR gains of up to 8 dB without degrading radar detection accuracy. Beyond waveform-level coordination, [9349171] studied uplink sensing under asynchronous transceivers and constrained antenna resources. Cross-antenna cross-correlation accurately estimates delays and Doppler shifts without strict synchronization, effectively resolving multi-target ambiguities by exploiting spatial correlation across receive antennas.
@@ III-B3 Topology of Multi-Cell Network
Figure 6: Topologies of Space-air-ground Integrated Network
@@
C-RAN, introduced around 2009, further advances the multi-cell paradigm by decoupling and centralizing BS functionalities. In LTE and 5G NR, a BS (gNB) can be functionally split into centralized units for higher-layer baseband processing, distributed units for lower-layer real-time tasks, and remote radio units (RRUs) for analog transmission and reception [3gpp_tr38801, TS38401]. To address full-duplex challenges, some RRUs are designated as dedicated receivers. When combined with massive MIMO and cooperative beamforming, C-RAN enables large-scale, coordinated wide-area sensing. Joint signal- and data-level fusion in C-RAN has been shown to improve both communication throughput and sensing accuracy [9737357]. In high-density networks, optimizing cluster size, user density, and spatial resource allocation to fully leverage the performance gains [10735119].
@@
The progression from distributed BS cooperation to heterogeneous macro–micro deployments and centralized C-RAN frameworks highlights the growing complexity of multi-cell ISAC. While these architectures offer extended coverage, spatial diversity, and improved sensing accuracy, they also introduce challenges in inter-cell synchronization, interference management, and coordination overhead. Advanced signal processing methods, such as compressed sensing and background subtraction, have been applied to recover perception parameters and mitigate multipath distortion [rahman2019framework]. A stochastic geometry model in [10735119] demonstrated that inter-cell coordination reduced sensing interference by up to 12% and improved SINR by nearly 5 dB. Moreover, joint beamforming and SINR optimization further improved worst-case radar SNR by 8% under communication constraints [10304081]. These findings underscore the need for unified cross-layer frameworks that can balance S&C demands in dynamic multi-cell environments.
@@ III-B4 Space-Air-Ground Integrated Network
The space–air–ground integrated network has been proposed to provide seamless connectivity and joint S&C services across multiple layers. As shown in Fig. 6, this architecture supports applications such as environmental monitoring, disaster response, and remote communications in areas where conventional infrastructure is limited. By integrating satellites, aerial platforms, and terrestrial networks, it not only extends coverage but also enables large-scale, real-time data acquisition and situational awareness across diverse geographic environments[yuan2025groundskyarchitecturesapplications].
@@
Despite its promise, SAGIN introduces unique technical hurdles. Spectrum sharing across segments is complicated by rapidly varying interference conditions, which necessitate collaborative spectrum sensing approaches, such as the distributed cognitive satellite–ground framework proposed in [hu2020energy]. Latency and real-time data processing pose additional challenges in high-mobility aerial scenarios, where distributed edge computing and predictive beamforming techniques [10745905] enable low-latency and mobility-resilient service continuity. Moreover, transmitting sensitive multi-segment data raises security and privacy concerns. Blockchain-based schemes have been proposed to guarantee tamper-proof, decentralized data integrity across space, air, and ground domains.
@@ III-C1 How Can Clock Asynchronization Be Effectively Mitigated in Distributed ISAC Networks
Clock asynchronization remains a fundamental limitation in large-scale ISAC deployments. Misaligned oscillators introduce timing offsets, carrier frequency shifts, and phase noise, thereby degrading range–Doppler estimation and symbol detection. Research has investigated mitigation through CSI ratio-based models [10678871], CRB-based performance analysis [9617146], and advanced estimators such as iterative cross-correlation [10273396], Kalman filtering [hamilton2008aces], and optimization-based offset estimation [10130780]. Practical implementations, such as the SPACE framework for joint synchronization and positioning and anchor-assisted uplink sensing [10637442], have integrated offset estimation and compensation. Despite these advances, scalable synchronization solutions for dynamic, multi-node ISAC networks remain limited, and trade-off analyses, such as those between synchronization accuracy and resource utilization, need further development.
@@ III-C2 How Should Sensing Architectures Be Designed to Balance Latency, Fusion Accuracy, and Scalability
Efficient sensing architecture design has been recognized as a critical enabler for the evolution of ISAC networks toward multi-cell and hierarchical deployments. At the 3GPP 6G workshop in March 2025 [6gworkshop], several candidate solutions were introduced, with a three-layer sensing framework emerging as one of the most intuitive options. This framework consists of three key components: (i) local sensing nodes, which include Remote Radio Units (RRUs), roadside units, and ISAC-enabled user equipment (UEs); (ii) intermediate aggregators, such as edge servers and distributed Baseband Units (BBUs); and (iii) the sensing function implemented within the core network. In this architecture, environmental signals are captured and pre-processed at the local nodes. The intermediate aggregators then collect and refine the measurements from multiple sources. Finally, the core network enforces system-wide requirements and carries out advanced data fusion and analytics. By distributing initial processing and partial fusion to these intermediate layers—rather than transmitting raw sensor data directly to the core—this hierarchical design reduces transmission latency, alleviates computational burdens on the sensing nodes, and mitigates the delays associated with fully centralized processing.
@@ IV-A From Edge Intelligence to Edge Perception
Modern smart environments are supported by billions of edge devices (e.g., smartphones, wearable sensors, and connected vehicles) that generate massive volumes of heterogeneous data. Traditional cloud-centric approaches, which offload computation to centralized servers, often incur high latency, excessive communication overhead, and elevated privacy risks [9815195]. Edge intelligence addresses these limitations by distributing computation across edge nodes, enabling localized data analysis and decision-making[LiTWC2020, Gartner]. Federated learning (FL), in particular, offers a privacy-preserving solution by training models locally while aggregating global insights at edge servers[Cao-ComM, 10478029].
@@ IV-A2 Edge Perception for Sensing Enhancement
Edge perception plays a central role in enhancing sensing capabilities through multi-modal data fusion at the network edge. By employing advanced AI algorithms, subtle patterns that elude conventional methods can be captured, thereby improving detection accuracy and situational awareness. For instance, in smart homes, sensor data is monitored to assess occupant behavior and environmental factors, which in turn facilitate the automatic optimization of temperature, humidity, and lighting to enhance comfort. In industrial IoT, real-time factory monitoring is utilized to support predictive maintenance and production scheduling, while wearable devices track patient vitals to enable prompt medical interventions. Beyond indoor and industrial settings, edge perception also plays a critical role in highly dynamic outdoor environments. In dense low-altitude operations, graph neural network–based methods enable UAVs to model local state information, extract multi-modal sensing features for conflict-risk awareness, and improve the effectiveness of autonomous avoidance decisions in large-scale swarms [liMultiscaleGraphEnhanced2025, liHierarchicalConflictResolution2025].
@@ IV-A3 Edge Perception for Communication Enhancement
In addition to sensing, edge perception strengthens communication performance by fusing diverse sensing modalities into the transmission process. Localized processing reduces end-to-end latency and alleviates computational pressure on centralized infrastructure. Multi-modal fusion improves communication reliability and resource management. In vehicular networks, radar, camera, and LiDAR inputs can be combined with RF signals to enhance interference mitigation, beamforming precision, and spectrum allocation. This synergy increases throughput and reduces latency. Beyond resource optimization, reinforcement learning and cross-attention architectures dynamically allocate spectrum resources, further improving adaptability and robustness in highly dynamic environments [shang2025energyefficientintelligentisacv2x].
@@ IV-A4 Interplay between ISAC and AI
This synergy has been validated across multiple application domains. In vehicular networks, a transformer-based predictive beamforming scheme that jointly exploited radar and CSI data improved beam selection accuracy and link throughput under mobility, achieving up to 4.2 × gains in low-speed scenarios [zhang2024transformer]. In indoor sensing, vision transformer–based models effectively processed Wi-Fi CSI data for human activity recognition (HAR). In particular, the class-attention in image transformers architecture achieved state-of-the-art classification accuracy on the University of Texas Human Activity Recognition (UT-HAR) dataset under constrained edge inference conditions [10477406]. Similarly, lightweight CSI image-based frameworks demonstrated high recognition accuracy while requiring only 2.1 MB of memory and delivering 18 frames per second throughput on Raspberry Pi-class hardware [10190332].
Beyond these applications, ISAC-enabled multimodal observation also strengthens multi-agent decision-making in highly dynamic environments. In UAV swarms, multi-agent deep reinforcement learning methods based on spatiotemporal graph reasoning and transformer-based sequence modeling enable efficient edge-node communication and coordination, significantly improving cooperative pursuit success probabilities against highly maneuverable targets under uncertain environments [yangSpatiotemporalGraphReasoning2025, yangTransformerbasedCooperativeUAV2025, yangCooperativePursuitevasionLow2025].
@@
This co-synergistic paradigm has been exemplified in [10621049], where an ISAC-guided over-the-air federated learning framework was introduced. By leveraging sensing data quality for dynamic client scheduling and power control, the framework improved convergence robustness and reduced training latency in bandwidth-limited environments[11143883]. To further enhance adaptability, [10608156] proposed a transfer learning–based edge model adaptation framework, which enabled fine-tuning of pretrained models using few-shot local samples. This design preserved model accuracy and privacy while significantly lowering communication overhead, demonstrating the potential of ISAC-assisted adaptive learning in dynamic wireless contexts.
@@
Energy efficiency and sensing cost further represent significant barriers in resource-constrained deployments. To alleviate these constraints, a closed-loop AI-in-the-loop framework was introduced in [cai2025aiintheloopsensingcommunicationjoint], where gradient importance sampling and stochastic gradient Langevin dynamics (SGLD)-based noise modeling were applied, achieving a 52% reduction in sensing cost and a 77% saving in communication energy. In passive sensing scenarios, a deep unfolding architecture named ISAC-NET was proposed in [10474422], alternating between channel reconstruction and target detection. This architecture yielded a 25.6% improvement in sensing accuracy and a 31.4% reduction in channel estimation error compared with conventional methods.
@@ IV-B1 The Shift towards Task-Oriented ISAC
Figure 10: The structure in paper[xu2024penetrative]
@@ IV-B2 Enabling AI Technologies
[li2021mmwave] High-pass filtering on range-Doppler data CNN Gesture recognition Accuracy >96%
[yadav2022tinyradar] Noise filtering and data normalization Posture detection Accuracy >93%
[salami2022teslarapture] Sparse point cloud processing CNN + BiLSTM + kNN Gesture recognition Accuracy >98%
@@
[Zhou2023BFPNetADL] Kalman filtering for motion state estimation Beam tracking in IoV systems Accuracy >92%, reduced beam misalignment by 30%
[grobelny2022mm] Fast fourier transform (FFT) for radar signal processing, noise filtering LSTM Hand gesture recognition for contactless interfaces (vending machines, ATMs, etc.) Accuracy >92%, reduced beam misalignment by 30%
@@
[signals3020017] Cell average constant false-alarm rate algorithm, data filtering Human activity recognition (boxing, jumping, squatting, etc.) Recognition accuracy >97%, 1% higher than using single data types
[hayashi2021radarnet] Miniature radar signal processing Gesture recognition Inference time <0.15 ms, Accuracy >90%
[zhang2024transformer] Echo signal processing, spatial filtering Transformer Predictive beamforming for vehicular ISAC systems Communication sum rate >3.7 dB compared to baseline, improved performance in high mobility
@@
[wang2022heartprint] Heartbeat-based radar sensing Supervised Learning Authentication (Heart rate data) Accuracy >97%
@@
Among recent breakthroughs, large language models (LLMs) have demonstrated strong potential in extracting knowledge from sensory data with minimal task-specific supervision. Typically structured into three modules, including perception, grounding, and alignment, LLMs enable multimodal reasoning and semantic abstraction[10638143]. For example, [wojcik2023beyond]introduced a framework where smartphone accelerometer signals were transformed into textual sequences for motion and environment inference, while electrocardiogram (ECG) signals were converted into text-based indicators such as heart rate, as shown in Fig. 10. Similarly, [xu2024penetrative] extended LLMs beyond textual processing to encode domain knowledge into cyber-physical systems. This signal-to-text paradigm highlights the semantic abstraction capabilities of LLMs in perception-centric ISAC networks.